# Community detection in network using Szegedy quantum walk

### v1

@@ Abstract
In a network, the vertices with similar characteristics construct communities. The vertices in a community are well-connected. Detecting the communities in a network is a challenging and important problem in the theory of complex networks. One approach to solve this problem uses the classical random walks on the graphs. In quantum computing, quantum walks are the quantum mechanical counterparts of classical random walks. In this article, we employ a variant of Szegedy’s quantum walk to develop a procedure for discovering the communities in networks. The limiting probability distribution of quantum walks assists us in determining the inclusion of a vertex in a community. We apply our procedure of community detection on a number of graphs and social networks, such as the relaxed caveman graph, $l$-partition graph, Karate club graph, dolphin’s social network, etc.
Keywords: Clustering, Community detection, Szegedy quantum walk, Complex networks and graphs.
@@ 1 Introduction
Networks are everywhere! Human beings build up networks among themselves for their progress and prosperity. They establish roads, rails, airlines, and the internet. They all form networks [1, 2]. Networks are inherent in nature. The human brain and protein-protein interaction network are two important examples of networks in nature. Due to the significance of networks, network science is an important branch of science and engineering [3]. In a network, elements of similar characteristics are connected. It distributes all the components in a network into communities. Detecting the communities in a given network is an important problem in network science [4, 5, 6, 7].
@@
Combinatorial graphs [8, 9] form the mathematical foundation of networks. We use the words graphs and networks synonymously throughout the article. A graph is a combination of objects and connection between them, where the objects are represented by the vertices and connection are represented by edges between the vertices. We mention the mathematical definition of graphs in section 2 of this article. The real-world networks display big inhomogeneities in the distribution of edges. We observe that the distribution of edges is not only globally but also locally inhomogeneous. In a real-world network, we find high concentrations of edges within special groups of vertices and low concentrations between these groups. These groups are called communities.
@@
Detecting the communities in a network is not easy. There are many community detection algorithms which are computationally NP-hard, or NP-complete [10, 11, 12]. The communities in graphs are often called the clusters. Therefore, the problem of community detection is also known as the graph clustering problem. Due to the existence of edges between the vertices, we prefer to use the word community instead of cluster. There are different methods of community detection discussed in the literature [13, 14, 15, 16, 17]. The random walks on graphs provide a crucial method to detect the communities [18, 19, 20, 21]. A random walker travels between the vertices following the connectivity structure of the graph. As the movement of the walker is random, at any time instance, there is a probability distribution for getting the walker at different vertices. Therefore, we get a sequence of probability distributions that converges to a limiting probability distribution under certain conditions. Usually, the inter-community edges are higher than the probability of the intra-community edges. In this way, the limiting probability distribution of random walks assists us in detecting the communities in graphs.
@@
In quantum computing, the quantum walks [22] are the quantize versions of classical random walks on the graphs [23, 24]. There are multiple proposals of quantum walks, such as discrete-time quantum walks [25], continuous-time quantum walks [26], open quantum walks [27], etc. The Szegedy quantum walk [25], is a variant of the discrete-time quantum walk. It converts the transition matrix of a classical random walk to a unitary matrix to govern the quantum walker on graphs. In addition we can define it on all the simple graph. Therefore, the Szegedy quantum walk is an efficient way to replace the idea of a classical random walk for developing a quantum computational algorithm.
@@
The graph clustering problem has recently been considered in the literature of quantum computing [28, 29, 30]. The quantum walks are also introduced for finding the clusters in graphs [31, 32]. Article [32] utilizes the discrete-time coined quantum walks to identify the communities in different real-world networks, such as Zachary’s karate club graph and the dolphin social networks and airline networks. The Fourier and Grover coins are used as quantum coin operator. It is demonstrated that the Fourier coin quantum walks localize within communities, which effectively reveal the modular structures of networks. In [31], the authors develop another graph clustering algorithm using discrete-time quantum random walks. Their method improves the clustering speed and accuracy. Article [30] discusses another efficient approach to clustering on directed and mixed graphs without using the idea of quantum walks.
@@
The motivation of this work is developed from our study $G$ the Szegedy quantum walk on the Barbell graphs. A barbell graph has two communities of vertices that are joined to a vertex by two edges and a vertex. The formal description will be mentioned later in this article. We observed that the limiting probability distribution of Szegedy quantum walk from a particular initial state has higher probability values for the inter-community edges. Therefore, using the Szegedy quantum walks, we can determine the inter community edges, whose removal generates the communities in the graph. This observation assists us in developing a method for community detection in networks.
@@
In this article, we present a new method for finding the communities in networks, which is applicable to all simple graphs with communities. There are multiple formalisms of Szegedy quantum walk, reported in literature [25]. For our constructions, we apply model discussed in [25]. To the best of our knowledge, this is the first research article where Szegedy quantum walk is applied for community detection. The quantum walk on graphs provides a limiting probability distribution on the edges of the graph. It assists us in determining a vertex to include in a community. For the numerical experiments, we apply our community detection techniques on relaxed caveman graphs, $l$-partition graphs, Zachary’s karate club graph, etc. The relaxed caveman graphs, $l$-partition graphs, are employed in the literature to examine different community detection methods [32, 5, 33]. Our result also matches with them. Due to the difference between the classical and quantum walks, we observe that the communities in Zachary’s karate club graph and the dolphin social network our communities are different from the communities derived in classical cases.
@@
The structure of this article is as follows: The primary concepts to go through this article, such as graphs, Szegedy quantum walks, etc., are discussed in section 2. In section 3, we describe our procedures of community detection using Szegedy’s quantum walk. The section 4 contains the communities in different graphs and social networks generated by our procedure. Then we conclude the article. There are two Appendix of this article. We consider a specific quantum state depending on the graphs for fast convergence of the probability distributions to the desired limiting probability distribution. In Appendix A, we discuss the choice of the particular quantum state. In Appendix B, we explain the convergence of the probability distribution.
@@ 2.1 A brief Overview of graph theory
A graph $G=(V(G),E(G))$ is a combinatorial structure, consists of a set of vertices $V(G)$ and a set of edges $E(G)\subseteq V(G)\times V(G)$ [8]. Throughout this article, $n$ and $m$ denote the number of vertices and edges, respectively, in a graph $G$, which are the finite natural numbers. If there is a direction on an edge we describe it as a directed edge. An undirected edge has no direction on it. An undirected edge between the vertices $u$ and $v$ is denoted by $(u,v)$. A loop is an edge joining a vertex with itself. An edge with an edge-weight is called an weighted edge. A graph without any directed edge, weighted edge, and loop is called a simple graph. Pictorially, we denote a vertex by a circle which may be marked or unmarked. A line joining two vertices $u$ and $v$ denotes an undirected edge $(u,v)$. A graph is said to be a complete graph if there is an edge between any two distinct vertices. The complete graph with $n$ vertices is denoted by $K_{n}$.
@@
In a simple graph $G$, two vertices $u$ and $v$ are said to adjacent if there is an edge $(u,v)\in E(G)$. Also, the edge $(u,v)$ is incident to the vertices $u$ and $v$. The number of incident edges to a vertex $v$ is called the degree of vertex $v$, which is denoted by $d_{G}(v)$. The neighborhood of a vertex $v$ in a simple graph $G$ is the set of all of its adjacent vertices $N(v)=\{v:(u,v)\in E(G)\}$. A path between the vertices $v_{0}$ and $v_{n}$ is a sequence of vertices and edges $v_{0},e_{1},v_{1},e_{2},v_{2},\dots e_{n},v_{n}$ where $e_{i}=(v_{i-1},v_{i})$ for $i=1,2,\dots n$, where no vertex is repeated. The number of edges in a path is called the length of a path. A graph is called connected if there is a path between any two vertices in it. In a connected graph $G$ there may be multiple paths between two vertices $u$ and $v$. The path with minimum length is called the shortest path between $u$ and $v$. A graph $G_{1}=(V(G_{1}),E(G_{1}))$ is said to be a subgraph of a graph $G=(V(G),E(G))$, if $V(G_{1})\subseteq V(G)$, and $E(G_{1})\subseteq E(G)$. A subgraph $G_{1}$ of $G$ is called an induced subgraph if $E(G_{1})=\{(u,v):(u,v)\in E(G)~\text{and}~u,v\in V(G_{1})\}$.
@@
The directed graphs play a crucial role in this article. To denote a directed edge between the vertices $u$ and $v$ we use the symbol $\overrightarrow{(u,v)}$. A directed graph is a graph having directed edges only. We denote a directed graph by $\overrightarrow{G}=(V(\overrightarrow{G}),E(\overrightarrow{G}))$. A directed edge $\overrightarrow{(u,v)}$ is represented by an arrow from $u$ to $v$. To denote two oppositely oriented edges $\overrightarrow{(u,v)}$ and $\overrightarrow{(v,u)}$ together, we use a straight line with two opposite arrows. Given a directed edge $\overrightarrow{e}=\overrightarrow{(u,v)}$, the vertex $u$ is said to be the origin of $\overrightarrow{e}$, and the vertex $v$ is the terminal of the edge $\overrightarrow{e}$. We denote $u=o(\overrightarrow{e})$, and $v=t(\overrightarrow{e})$. We say that $\overrightarrow{e}$ is an incoming edge to the vertex $v$. Also, $\overrightarrow{e}$ is an outgoing edge from the vertex $u$. The indegree of a vertex $u$ is the number of incoming edges to the vertex $v$, which we denote by $d^{(i)}_{\overrightarrow{G}}(v)$. Also, the outdegree of a vertex $u$ is the number of outgoing edges from $u$, which is denoted by $d^{(o)}_{\overrightarrow{G}}(v)$. The inverse of a directed edge $\overrightarrow{e}=\overrightarrow{(u,v)}$, is a directed edge $\overrightarrow{e}^{-1}=\overrightarrow{(v,u)}$. Therefore, $o(\overrightarrow{e})=t(\overrightarrow{e}^{-1})$ and $t(\overrightarrow{e})=o(\overrightarrow{e}^{-1})$.
@@
In this paper, we consider the connected, undirected graphs. A community in a graph is a set of vertices with the similar properties. In this article, we emphasize on the connectivity properties of the vertices. We assume that a vertex belongs to at most one community. Therefore, the set of communities generated by our procedure is crisp. The set of communities $\mathcal{C}=\{C_{1},C_{2},\dots,C_{k}\}$ in a graph is a partition on $V(G)$ where $C_{i}$ are the communities for $i=1,2,\dots k$. Therefore, $\cup_{i=1}^{n}C_{i}=V(G)$ and $C_{i}\cap C_{j}=\emptyset$ for $i\neq j$. We say $\mathcal{C}$ is trivial if either $k=1$, or $C_{i}$ is a singleton set for all $i$. We identify a community $C_{i}$ with the induced subgraph $G[C_{i}]=(C_{i},E(C_{i}))$ of the graph $G$ for every community $C_{i}$, where $E(C_{i}):=\{(v,w)\in E:v,w\in C_{i}\}$. The $E(\mathcal{C})=\cup_{i=1}^{k}E(C_{i})$ is the set of intra-community edges of the graph $G$ induced by the community detection $\mathcal{C}$. The number of edges in $E(\mathcal{C})$ is denoted by $m(\mathcal{C})$. Also, $E(G)-E(C)=\{(u,v):u\in C_{i},\ v\in C_{j},\ i\neq j\}$ is the set of inter-community edges of $G$ induced by the community detection $\mathcal{C}$. [34]. The number of intra-community edges is more than the number of inter-community edges.
@@ 2.2 An introduction to quantum walks
Let $G=(V(G),E(G))$ be a connected simple graph. We assign two opposite orientations on every undirected edge $(u,v)\in E(G)$. Therefore, from every undirected edge $(u,v)$ in $G$, we construct two directed edges $\overrightarrow{e}=\overrightarrow{(u,v)}$ and $\overrightarrow{e^{-1}}=\overrightarrow{(v,u)}$. The new directed graph is $\overrightarrow{G}$. Therefore, if $G$ has $m$ edges, then $\overrightarrow{G}$ has $2m$ directed edges. Let the edges $\overrightarrow{e_{1}},\overrightarrow{e_{2}},\dots,\overrightarrow{e_{m}},\overrightarrow{e_{1}^{-1}},\overrightarrow{e_{2}^{-1}},\dots,\overrightarrow{e_{m}^{-1}}$ are marked by the integers $0,1,2,\dots(2m-1)$. The construction of $\overrightarrow{G}$ from $G$ indicates that $d_{G}(v)=d^{(o)}_{\overrightarrow{G}}(v)=d^{(i)}_{\overrightarrow{G}}(v)$, where $d_{G}(v)$ is the degree of $v$ in $G$, as well as $d^{(o)}_{\overrightarrow{G}}(v),d^{(i)}_{\overrightarrow{G}}(v)$ are the out-degree and in-degree of $v$ in $\overrightarrow{G}$, respectively.
Let a walker is moving on the directed graph $\overrightarrow{G}$ following the graph structure. At a particular time $t$, let the walker arrives at a vertex $v$. In the next step, the walker will move to one of the adjacent vertices of $v$. Hence, the walker must follow one of the outgoing edges $\overrightarrow{e}$ from $v$. Recall that, if out-degree of $v$ is $d_{v}^{(o)}$, then there are $d_{v}^{(o)}$ out-going edges from $v$. The walker will choose one of them with probability $p(\overrightarrow{e})$. Therefore, we can define a function $p:E(\overrightarrow{G})\rightarrow(0,1]$ satisfying
$\sum_{\overrightarrow{e}:o(\overrightarrow{e})=v}p(\overrightarrow{e})=1,$ (1)
for each vertex $v\in V$. If we assume that choosing an outgoing edge from $v$ is equiprobable, then $p(\overrightarrow{e})=\frac{1}{d^{(o)}_{v}}=\frac{1}{d_{v}}$, where $v=o(\overrightarrow{e})$.
Now, we are in a position to define the Szegedy quantum walk. We represent the state of a walker by a quantum state vector. Corresponding to the directed graph $\overrightarrow{G}$ with $2m$ edges, we consider a Hilbert space $\mathcal{H}^{2m}$ of dimension $2m$. The basis of $\mathcal{H}^{2m}$ is $\left\{|0\rangle,|1\rangle,|2\rangle,\dots,|2m-1\rangle\right\}$, where
[…]
@@
Here, $\delta_{(\overrightarrow{e},\overrightarrow{f})}$ is the Kronecker delta, which is defined by
$\delta_{(\overrightarrow{e},\overrightarrow{f^{-1}})}=\begin{cases}1&\text{if}~\overrightarrow{e}=\overrightarrow{f};\\ 0&\text{otherwise}.\end{cases}$ (4)
@@
The procedure of clustering depends on the selection of the initial state. A different choice of initial state may lead us to another limiting probability distribution, which may not be useful for us. One such example is presented in Appendix A
3 Community detection on graphs
[…]

### v2

@@ Abstract
In a network, the vertices with similar characteristics construct communities. The vertices in a community are well-connected. Detecting the communities in a network is a challenging and important problem in the theory of complex networks. One approach to solving this problem uses the classical random walks on graphs. In quantum computing, quantum walks are the quantum mechanical counterparts of classical random walks. In this article, we employ a variant of Szegedy’s quantum walk to develop a procedure for discovering the communities in networks. The limiting probability distribution of quantum walks assists us in determining the inclusion of a vertex in a community. We apply our community detection procedure to a variety of graphs and social networks, including the relaxed caveman graph, $l$-partition graph, Karate club graph, and the dolphin’s social network, among others.
Keywords: Clustering, Community detection, Szegedy quantum walk, Complex networks.
@@ 1 Introduction
Networks are everywhere! Human beings build up networks among themselves for their progress and prosperity. They establish roads, rails, airlines, and the internet. They all are different networks [1, 2]. Networks are inherent in nature. The human brain and protein-protein interaction network are two important examples of networks in nature [3]. In a network, elements of similar characteristics are well-connected. It distributes all the elements in a network into communities. Detecting the communities in a given network is an important problem in network science [4, 5, 6, 7].
@@
Combinatorial graphs [8, 9] form the mathematical foundation of networks. We use the words graphs and networks synonymously throughout the article. A graph is a combination of objects and connection between them, where the objects are represented by the vertices and connection are represented by edges between them. The real-world networks display big inhomogeneities in the distribution of edges. The concentrations of edges within special groups of vertices is high. We call them communities, or clusters. Due to the existence of connections between the vertices, we prefer to use the word community instead of cluster.
@@
Detecting the communities in a network is not an easy task. There are different methods of community detection discussed in the literature [10, 11, 12, 13, 14, 15]. Many of the community detection algorithms are computationally NP-hard, or NP-complete [16, 17, 18]. The random walks on graphs provide a crucial method to detect the communities [19, 20, 21, 22]. A random walker travels between the vertices following the connectivity structure of the graph. As the movement of the walker is random, at any time-instance, there is a probability distribution for getting the walker at different vertices. Therefore, we get a sequence of probability distributions that converges to a limiting probability distribution under certain conditions. Usually, the inter-community edges are higher than the probability of the intra-community edges. In this way, the limiting probability distribution of random walks assists us in detecting the communities.
@@
In quantum computing, the quantum walks [23] are the quantize versions of classical random walks on the graphs [24, 25]. The quantum walks are applied in quantum communication [26, 27], and computation [28, 29, 30]. There are multiple proposals of quantum walks, such as discrete-time quantum walks [31], continuous-time quantum walks [32], open quantum walks [33, 34], etc. The Szegedy quantum walks [31] converts the transition matrix of a classical random walk to a unitary matrix to govern the quantum walker on graphs. In addition we can define it on all the simple graph. Therefore, the Szegedy quantum walk is an efficient way to replace the idea of random walk for developing a quantum algorithm.
@@
The graph clustering problem has recently been considered in the literature of quantum computing [35, 36, 37]. The quantum walks are also introduced for finding the clusters in graphs [38, 39]. Article [39] utilizes the discrete-time coined quantum walks to identify the communities in different real-world networks. The Fourier and Grover coins are used as the quantum coin operators. It is demonstrated that the Fourier coin quantum walks localize within communities, which effectively reveal the modular structures of networks. Article [38] improves the clustering speed and accuracy. Article [37] discusses another efficient approach to clustering on directed and mixed graphs without using quantum walks.
@@
The motivation of this work comes from the Szegedy quantum walk on the Barbell graphs. A barbell graph has two communities of vertices that are joined to a central vertex by two inter-community edges. Ref. 2(a). We observed that the limiting probability distribution of Szegedy quantum walk from a particular initial state has higher probability values for the inter-community edges. Their removal generates the communities in the graph. This observation assists us in developing a method for community detection in networks.
@@
In this article, we present a new method for finding the communities in networks, which is applicable to all simple graphs with communities. For our constructions, we apply the model of Szegedy quantum walks discussed in [31]. To the best of our knowledge, this is the first research article where Szegedy quantum walk is applied for community detection. The quantum walk on graphs provides a limiting probability distribution on the edges of the graph. It assists us in determining a vertex to include in a community. For the numerical experiments, we apply our community detection techniques on different networks which are already used in literature [39, 5, 40]. Our results also matches with them. Due to the difference between the random walks and quantum walks, we observe that the communities are different in some cases.
@@
The structure of this article is as follows: The primary concepts are discussed in section 2. In section 3, we describe our procedures of community detection. The section 4 contains the communities in different networks generated by our procedure. Then we conclude the article. In Appendix we discuss the choice of our initial quantum state, and the convergence of the probability distribution.
@@ 2.1 A brief Overview of graph theory
A graph $G=(V(G),E(G))$ consists of a set of vertices $V(G)$ and a set of edges $E(G)\subseteq V(G)\times V(G)$ [8]. Throughout this article, $n$ and $m$ denote the number of vertices and edges, respectively. An undirected edge between the vertices $u$ and $v$ is denoted by $(u,v)$. A loop is an edge joining a vertex with itself. A graph with the undirected edges, and without loops is called a simple graph. We do not consider weighted graphs, in this article. A graph with $n$ vertices is said to be a complete graph, $K_{n}$ if there is an edge between any two distinct vertices.
@@
In a simple graph $G$, two vertices $u$ and $v$ are said to be adjacent if there is an edge $(u,v)$. Also, the edge $(u,v)$ is incident to the vertices $u$ and $v$. The number of incident edges to a vertex $v$ is called the degree of $v$, which is denoted by $d_{G}(v)$. The neighborhood of a vertex $v$ in $G$ is the set $N(v)=\{v:(u,v)\in E(G)\}$. A path between the vertices $v_{0}$ and $v_{n}$ is a sequence of vertices and edges $v_{0},e_{1},v_{1},e_{2},v_{2},\dots e_{n},v_{n}$ where $e_{i}=(v_{i-1},v_{i})$ for $i=1,2,\dots n$, where no vertex is repeated. The number of edges in a path is called the length of a path. A graph is called connected if there is a path between any two vertices. The path between $u$ and $v$ with minimum length is called the shortest path. A graph $G_{1}=(V(G_{1}),E(G_{1}))$ is said to be a subgraph of a graph $G=(V(G),E(G))$, if $V(G_{1})\subseteq V(G)$, and $E(G_{1})\subseteq E(G)$. A subgraph $G_{1}$ of $G$ is called an induced subgraph if $E(G_{1})=\{(u,v):(u,v)\in E(G)~\text{and}~u,v\in V(G_{1})\}$.
@@
A directed edge from $u$ to $v$ is denoted by $\overrightarrow{(u,v)}$. All the edges in a directed graph $\overrightarrow{G}=(V(\overrightarrow{G}),E(\overrightarrow{G}))$ is directed. Given a directed edge $\overrightarrow{e}=\overrightarrow{(u,v)}$, the vertices $u=o(\overrightarrow{e})$, and $v=t(\overrightarrow{e})$ are said to be the origin and the terminal of $\overrightarrow{e}$, respectively. The indegree of a vertex $v$ is denoted by $d^{(i)}_{\overrightarrow{G}}(v)$, which is the number of edges $\overrightarrow{e}$ such that $v=t(\overrightarrow{e})$. Also, the outdegree of a vertex $v$ is denoted by $d^{(o)}_{\overrightarrow{G}}(v)$, which is the number of edges $\overrightarrow{e}$ such that $v=o(\overrightarrow{e})$. The inverse of a directed edge $\overrightarrow{e}=\overrightarrow{(u,v)}$, is $\overrightarrow{e}^{-1}=\overrightarrow{(v,u)}$. Therefore, $o(\overrightarrow{e})=t(\overrightarrow{e}^{-1})$ and $t(\overrightarrow{e})=o(\overrightarrow{e}^{-1})$.
@@
In this article, we assume that a vertex belongs to at most one community. The set of communities $\mathcal{C}=\{C_{1},C_{2},\dots,C_{k}\}$ in a graph is a partition on $V(G)$, such that, $\cup_{i=1}^{n}C_{i}=V(G)$ and $C_{i}\cap C_{j}=\emptyset$ for $i\neq j$. We say that $\mathcal{C}$ is trivial if either $k=1$, or $C_{i}$ is a singleton set for all $i$. We identify a community $C_{i}$ with the induced subgraph $G[C_{i}]=(C_{i},E(C_{i}))$ of the graph $G$ for every community $C_{i}$, where $E(C_{i}):=\{(v,w)\in E:v,w\in C_{i}\}$. The $E(\mathcal{C})=\cup_{i=1}^{k}E(C_{i})$ is the set of intra-community edges induced by $\mathcal{C}$. The number of edges in $E(\mathcal{C})$ is denoted by $m(\mathcal{C})$. Also, $E(G)-E(C)=\{(u,v):u\in C_{i},\ v\in C_{j},\ i\neq j\}$ is the set of inter-community edges of $G$ induced by $\mathcal{C}$. The number of intra-community edges is more than the number of inter-community edges.
@@ 2.2 An introduction to quantum walks
From a connected simple graph $G=(V(G),E(G))$ we construct a directed graph $\overrightarrow{G}$ by replacing every undirected edge $(u,v)$ by two directed edges $\overrightarrow{e}=\overrightarrow{(u,v)}$ and $\overrightarrow{e^{-1}}=\overrightarrow{(v,u)}$. Let the edges $\overrightarrow{e_{1}},\overrightarrow{e_{2}},\dots,\overrightarrow{e_{m}},\overrightarrow{e_{1}^{-1}},\overrightarrow{e_{2}^{-1}},\dots,\overrightarrow{e_{m}^{-1}}$ in $\overrightarrow{G}$ are marked by the integers $0,1,2,\dots(2m-1)$. The construction of $\overrightarrow{G}$ from $G$ indicates that $d_{G}(v)=d^{(o)}_{\overrightarrow{G}}(v)=d^{(i)}_{\overrightarrow{G}}(v)$.
At time $t$, let the walker is at a vertex $v$. At time $(t+1)$, following one of the $d_{v}^{(o)}$ outgoing edges it reaches to a neighboring vertex. The walker chooses one of the out-going edges with probability $p(\overrightarrow{e})=\frac{1}{d^{(o)}_{v}}=\frac{1}{d_{v}}$, where $v=o(\overrightarrow{e})$. Therefore, $\sum_{\overrightarrow{e}:o(\overrightarrow{e})=v}p(\overrightarrow{e})=1$, for each vertex $v\in V$.
@@
Here, $\delta_{(\overrightarrow{e},\overrightarrow{f^{-1}})}$ is the Kronecker delta, which is defined by
$\delta_{(\overrightarrow{e},\overrightarrow{f^{-1}})}=\begin{cases}1&\text{if}~\overrightarrow{e}=\overrightarrow{f^{-1}};\\ 0&\text{otherwise}.\end{cases}$ (2)
@@
The procedure of community detection depends on the selection of the initial state, which is further explored in Appendix A. At $t\textsuperscript{th}$ step of quantum walk for $t=0,1,2,\dots$, the state vector of the walker is